The current in a circuit with an external resistance of 3.75 ohm is 0.5 . when a resistance of 1ohm is introduce into the circuit the current becomes 0.4 .Then what will be the emf of the power source ? Please explain .

Grade:12

1 Answers

Arun
5 years ago
E=2V\implies E=0.4(3.75+1+r)=0.4(4.75+r)E=i(R+R_1+r)is added to the circuit,1\OmegaWhen external resistance of\implies E=0.5(3.75+r), respectively.When external resistance of3.75 omega is connected,randEet the EMF and internal resistance of the battery be